A number is selected randomly from a container containing all the integers from
10 to 50 Find
a) P (even | greater than 40)
b) P (greater than 40 | even)
c) P (prime | between 20 and 40)

Answere this with solution .
1
Expert's answer
2019-07-24T06:05:53-0400

a) P (even | greater than 40) = P (even AND greater than 40) / P(greater than 40) = 5/41 * 41/10 = 0.5

b) P (greater than 40 | even) = P (greater than 40 AND even) / P(even) = 5/41 * 41/21 = 5/21

c) P (prime | between 20 and 40) = P (prime AND between 20 and 40) / P(between 20 and 40) = 4/41 * 41/19 = 4/19